Jenkins, Willis, and Trent invested $268,000, $469,000, and $603,000, respectively, in a partnership. During its first year, the firm recorded profit of $651,000.

Required:
Prepare entries to close the firm’s Income Summary account as of December 31 and to allocate the profit to the partners under each of the following assumptions:

a.  The partners did not produce any special agreement on the method of distributing profits.

  • Record to close income summary account

b. The partners agreed to share profit and losses in the ratio of their beginning investments.

  • Record to close income summary account.


    c.   The partners agreed to share profit by providing annual salary allowances of $127,000 to Jenkins, $137,000 to Willis, and $72,000 to Trent; allowing 15% interest on the partners’ beginning investments; and sharing the remainder equally.
  • Record to close income summary account.
Answer #1

Given ,capitals

Jenkins = 268000

willis =469000

Trent = 603000

a) if there is no special agreement to share profits than distribute profits to partners in equal equal ratio.

sharing profit =651000/3

=217000 to each partner

partnership profit 651000

Jenkins capital 217000

Willis capital 217000

Trent capital 217000

(being capital distributed to partners in equal ratio)

b) profits and losses sharing ratio = partner capital / total capital

jenkins ratio =268000 / 1340000 = 20%

jenkins profit = 651000* 20% =103200

willis ratio =469000 /1340000 =35%

willis profit =651000 * 35% =227850

trent ratio = 603000 / 1340000 =45%

trent profit = 651000 * 45% = 292950

partnership profit 651000

jenkins capital 103200

willis capital 227850

trent capital 292950

(being profit distributed in capital sharing ratio)

c)  share of partner = salary + interest 15% on capital + remainder share equally

jenkins = 127000 + 40200 + 38000

=205200

willis =137000 + 70350 + 38000

=245350

trent = 72000 + 90450 + 38000

=200450

partnership profit 651000

jenkins capital 205200

willis capital 245350

trent capital 200450

( being profit was distributed to partners)